# Negate or Embrace: On How Misalignment Shapes Multimodal Representation Learning

### v3

@@ 5.2 MPI3D-Complex Dataset
The MPI3D-Complex dataset (Gondal et al., 2019) contains real-world images annotated with seven mutually independent, discrete ground-truth factors: object color (color), shape (shape), size (size), camera height (cam.), background color (back.), horizontal position (hori.), and vertical position (vert.). We treat hori. and vert. as image-specific factors, and the remaining five as semantic variables. See Figure 5 for example images from the MPI3D-Complex dataset.
@@
Training is performed using the loss $\mathcal{L}_{\text{MMCL}}$ defined in Eq. 1, following the formulation in (Daunhawer et al., 2022). Performance is evaluated using the average Matthews Correlation Coefficient (MCC) across three random seeds, based on the prediction of all image latent variables from the learned image and text representations using linear and nonlinear MLP decoders, respectively. Full details on dataset attributes, text generation, bias configurations, and architectures are provided in § D.1.
@@ Results.
Figure 6 presents the nonlinear MCC results with learned representations. The findings demonstrate that, even with discrete latent variables, misaligned semantic variables across modalities, whether due to selection or perturbation biases, are systematically excluded from the representations (MCC $=0$). In contrast, unbiased semantics are well recovered, with MCC scores predominantly approaching 1 and all values $\geq 0.8$, reinforcing our theoretical findings. Further investigations into MCC using linear classifiers and ablation studies on encoder dimensionality are provided in § D.2.
@@ 5.3 Causal3DIdent Dataset
We further conduct experiments on the Causal3DIdent dataset (Zimmermann et al., 2021; Von Kügelgen et al., 2021; Daunhawer et al., 2022; Yao et al., 2023), a semi-synthetic dataset allowing explicitly defined causal relations among latent variables. Images are generated from 10 latent variables: 3 discrete (shape, x_pos, y_pos) and 7 continuous (color, s_pos, s_color, b_color, alpha, beta, gamma). We treat the rotation angles (alpha, beta, gamma) as image-specific, and the remaining as semantic variables following a causal graph in Figure 15. See Figure 7 for example images from the Causal3DIdent dataset.
@@ Results.
Figure 8 presents the prediction performance of a nonlinear MLP classifier or regressor trained on learned image representations. We observe that unbiased semantic variables, whether continuous (e.g., s_pos) or discrete (e.g., shape, x_pos, y_pos), are reliably captured across all settings, with predictive performance approaching perfect ($R^{2}\approx 1$). For semantic variables that are continuous in the image modality but discretized in the text modality, prediction performance shows some degradation (e.g., s_color in selection setting 6 or perturbation setting 2), yet still achieves relatively high $R^{2}$ scores. Image-specific variables are consistently excluded from the learned representations, as indicated by $R^{2}=0$.
@@ 6 Related Work
Recent work has sought to formalize the theoretical foundations of multimodal and multi-view representation learning, particularly under contrastive objectives. Wang et al. (Wang and Isola, 2020) decompose the InfoNCE loss (Oord et al., 2018) into an alignment term, which pulls positive pairs together, and a uniformity term, which encourages dispersion over a hypersphere—laying the groundwork for subsequent analysis. Zimmermann et al. (Zimmermann et al., 2021) show that contrastive objectives can invert the data-generating process, while Liu et al. (Liu et al., 2024d) extend this result to multimodal settings. However, these approaches often rely on strong assumptions about latent distributions or manifold structures, which limits their practical applicability. A complementary line of work, such as (Von Kügelgen et al., 2021), demonstrates that contrastive learning with data augmentation can recover shared content without requiring such strong assumptions. This has been extended to MMCL by (Lyu et al., 2022; Daunhawer et al., 2022), and Yao et al. (Yao et al., 2023) further investigate identifiability under partial observability in multi-view settings. Distinct from prior work, we do not assume fixed content–style decompositions, causal directions, or perfectly aligned pairs. Instead, we analyze MMCL with image–text pairs under misalignment and systematically examine its impact on representation learning.
@@
Multimodal contrastive learning (MMCL) has achieved significant empirical success, particularly in aligning visual and textual modalities using models such as CLIP (Radford et al., 2021) and ALIGN (Jia et al., 2021). These successes are partly attributed to the use of massive training corpora, e.g., LAION-5B (Schuhmann et al., 2022), which are substantially larger than those used for vision-only foundation models (He et al., 2022; Oquab et al., 2024). However, real-world multimodal datasets are often imperfectly aligned and noisy (Miech et al., 2019). Existing empirical methods (Alayrac et al., 2020; Lavoie et al., 2024) typically treat such misalignment as label noise, employing strategies such as multiple-instance learning or dataset refinement (Alayrac et al., 2020) to mitigate its impact. While some recent work suggests that contrastive models are robust to certain forms of structured misalignment (Nakada et al., 2023), others propose augmenting text to simulate semantic variation in visual content (Cai et al., 2025). Our work suggests that misalignment can act as either a barrier or a bridge, depending on the application. Unlike (Nakada et al., 2023), which assumes linear representations without modeling the generative process, our analysis is grounded in a realistic latent variable model that provides a deeper understanding of misalignment from a data-generating perspective.
@@
Identifiability analysis addresses the fundamental question of whether the learning process can uniquely recover the latent generative structure or distribution underlying the observed data. This problem has been extensively studied in the context of nonlinear independent component analysis (ICA) (Hyvärinen and Pajunen, 1999; Locatello et al., 2019; Khemakhem et al., 2020; Sorrenson et al., 2020) and causal representation learning (Liu et al., 2022, 2024c; Zhang et al., 2024a; Yao et al., 2025; Liu et al., 2025). In practice, full identifiability—typically up to permutation—is rarely achievable without strong assumptions. Consequently, recent works have focused on partial identifiability (Gu and Xu, 2020; Kong et al., 2023; Gao et al., 2025) or relaxed equivalence classes, such as identifiability up to linear transformations (Zimmermann et al., 2021; Liu et al., 2024d) or up to group-wise/block-wise indeterminacy (Von Kügelgen et al., 2021; Yao et al., 2023), which can offer sufficient guarantees for specific tasks or settings. In the context of multimodal representation learning, several recent studies have explored identifiability results (Lyu et al., 2022; Daunhawer et al., 2022; Liu et al., 2024d), but largely neglect the presence of systematic misalignment. In contrast, our work explicitly models misalignment and adopts a block-identifiability definition to characterize the extent to which semantic factors can be recovered up to an invertible mapping.
@@
Invariant representation learning (IRL) seeks to learn representations that remain robust under distributional shifts between environments (Arjovsky et al., 2019; Eastwood et al., 2022; Gao et al., 2025), particularly in settings where empirical risk minimization (ERM) (Vapnik, 1991) fails to generalize out of distribution. In the absence of such variation, ERM is sufficient for in-distribution prediction, rendering the objective of IRL ill-posed. From a causal perspective, learning invariant representations—or more ambitiously, invariant mechanisms—requires variability in the non-invariant factors of the data. Such variability can be introduced through interventional data (Lippe et al., 2022; Lachapelle et al., 2022), exchangeability assumptions (Reizinger et al., 2025), or the use of auxiliary variables such as domain indices (Zhang et al., 2024a; Liu et al., 2024b). However, direct interventions on latent variables are typically infeasible in real-world data, and auxiliary variable methods often require access to a large number of diverse environments to ensure identifiability—an assumption that is rarely satisfied in practice. Our work offers an alternative approach: by leveraging the inherent flexibility of text supervision, we demonstrate that manipulating biases—specifically through selective omission or semantic perturbation in text—can serve as a controllable proxy for environmental variation.
@@ 7 Conclusion and Discussion
In this work, we present a formal analysis of misalignment in MMCL, examining its impact on learned representations. We demonstrate that contrastive multimodal encoders retain only the unbiased shared semantic variables, systematically discarding misaligned latent variables. When image-text pairs exhibit selection or perturbation biases, the joint embedding prioritizes consistent content, while omitting altered or missing aspects. This trade-off is fundamental: perfectly aligned text captions preserve rich semantic detail, whereas selective or biased text can enhance domain invariance by filtering out distribution-sensitive factors. Our experiments, conducted across simulations and image-text datasets, empirically validate these theoretical findings. These insights underscore the need for multimodal learning frameworks that either mitigate misalignment or leverage beneficial biases to improve representation learning in real-world settings. A more detailed discussion of these implications is provided in App. G.
@@ Appendix A Notation and Terminology
$\mathbb{I}_{\mathbf{s}}$ Index set of semantic variables: $\{1,\dots,n_{s}\}$
@@
$\mathcal{L}_{\text{MMCL}}(\mathbf{f}_{x},\mathbf{f}_{t})$ Symmetric InfoNCE loss for MMCL (Eq. 1)
$\mathcal{L}_{\text{SymAlignMaxEnt}}(\mathbf{f}_{x},\mathbf{f}_{t})$ Alignment and entropy maximization loss (Eq. 2)
Notations for Misalignment
$\theta$ Selection bias, an integer realization in the range $\{1,\dots,2^{n_{s}}-1\}$
@@
$\rho$ Perturbation bias, an integer realization in the range $\{1,\dots,2^{|\mathbb{I}_{\theta}|}-1\}$
@@
$\mathbf{s}_{\mathbb{I}^{c}_{\theta}}$ Omitted latent semantic variables for generating text
@@
$\mathbf{z}_{x}$ Combined latent variable for images: $(\mathbf{s},\mathbf{m}_{x})$
$\mathbf{z}_{t^{\scriptscriptstyle(\theta)}}$ Combined latent variable for text view: $(\mathbf{s}_{\mathbb{I}_{\theta}},\mathbf{m}_{t})$
@@ B.1 Lemmas
Lemma B.1 (Global Minimum of $\mathcal{L}_{\text{SymAlignMaxEnt}}$).
Under the assumptions of Thm. 4.1, the global minimum of
$\mathcal{L}_{\text{SymAlignMaxEnt}}(\mathbf{f}_{x},\mathbf{f}_{t})=\mathop{\mathbb{E}}_{(\mathbf{x},\mathbf{t}^{\scriptscriptstyle(\theta)})\sim p_{\mathbf{x},\mathbf{t}^{\scriptscriptstyle(\theta)}}}\left[\|\mathbf{f}_{x}(\mathbf{x})-\mathbf{f}_{t}(\mathbf{t}^{\scriptscriptstyle(\theta)})\|_{2}\right]-\frac{1}{2}\Bigl{(}H\bigl{(}\mathbf{f}_{x}(\mathbf{x})\bigr{)}+H\bigl{(}\mathbf{f}_{t}(\mathbf{t}^{\scriptscriptstyle(\theta)})\bigr{)}\Bigr{)},$ (6)
@@
  • The operator $(\cdot)_{\mathbb{I}^{c}_{\rho}}$ extracts the components corresponding to the preserved semantic variables (i.e., unaffected by the selection bias nor the perturbation bias), with $n=\left|\mathbb{I}^{c}_{\rho}\right|$ being their dimensionality.
  • $\mathbf{d}=(d_{1},\dots,d_{n})$ is defined via the Darmois construction (Darmois, 1951; Hyvärinen and Pajunen, 1999; Von Kügelgen et al., 2021), where for each $i\in[n]$ (we abbreviate $\{1,\cdots,n\}$ as $[n]$ for any integer $n>1$ for simplicity),
@@
We prove that the candidate functions $\mathbf{f}_{x}^{*}$ and $\mathbf{f}_{t}^{*}$ in Equations 7 and 8 yield $\mathcal{L}_{\text{SymAlignMaxEnt}}(\mathbf{f}_{x}^{*},\mathbf{f}_{t}^{*})=0$. Substituting these candidate functions into the loss in Eq. 6, we have
$\mathcal{L}_{\text{SymAlignMaxEnt}}(\mathbf{f}_{x}^{*},\mathbf{f}_{t}^{*})$ $=\mathop{\mathbb{E}}_{(\mathbf{x},\mathbf{t}^{\scriptscriptstyle(\theta)})\sim p_{\mathbf{x},\mathbf{t}^{\scriptscriptstyle(\theta)}}}\Bigl{[}\|\mathbf{f}_{x}^{*}(\mathbf{x})-\mathbf{f}_{t}^{*}(\mathbf{t}^{\scriptscriptstyle(\theta)})\|_{2}\Bigr{]}$
@@
$\mathcal{L}_{\text{SymAlignMaxEnt}}(\mathbf{f}_{x}^{*},\mathbf{f}_{t}^{*})$ $=\mathop{\mathbb{E}}_{(\mathbf{z}_{x},\mathbf{z}_{t^{\scriptscriptstyle(\theta)}})\sim p_{\mathbf{z}_{x},\mathbf{z}_{t^{\scriptscriptstyle(\theta)}}}}\Bigl{[}\|\mathbf{d}(\mathbf{s}_{\mathbb{I}^{c}_{\rho}})-\mathbf{d}(\tilde{\mathbf{s}}_{\mathbb{I}^{c}_{\rho}})\|_{2}\Bigr{]}$
@@
$\tilde{\mathbf{s}}_{\mathbb{I}^{c}_{\rho}}=\mathbf{s}_{\mathbb{I}^{c}_{\rho}}\quad\text{almost surely (a.s.)}\quad\forall\,\mathbf{s}\sim p_{\mathbf{s}},\tilde{\mathbf{s}}_{\mathbb{I}_{\theta}}\sim p_{\tilde{\mathbf{s}}_{\mathbb{I}_{\theta}}\mid\mathbf{s}_{\mathbb{I}_{\theta}},\rho},$
@@
Further, by the properties of the Darmois construction (Darmois, 1951), the mapping $\mathbf{d}$ transforms $\mathbf{s}_{\mathbb{I}^{c}_{\rho}}$ into a uniform distribution over $(0,1)^{n}$ (with $n=|\mathbb{I}^{c}_{\rho}|$). Since the uniform distribution is the unique maximum entropy (i.e., zero) distribution on a bounded domain (under no further moment constraints) (Jaynes, 1982; Cover, 2006), the entropy terms in the loss are maximized. In the formulation of $\mathcal{L}_{\text{SymAlignMaxEnt}}$, this maximal entropy precisely cancels any potential reduction in the loss, ensuring that
$\mathcal{L}_{\text{SymAlignMaxEnt}}(\mathbf{f}_{x}^{*},\mathbf{f}_{t}^{*})=0.$
Therefore, the global minimum of $\mathcal{L}_{\text{SymAlignMaxEnt}}$ is achieved at 0 by the given function pairs $\mathbf{f}_{x}^{*}$ and $\mathbf{f}_{t}^{*}$, completing the proof. ∎
@@
Let $\mathbf{h}:\mathcal{U}\to\mathcal{V}$ be a smooth map between simply connected, open $\mathcal{C}^{1}$ manifolds $\mathcal{U},\mathcal{V}\subseteq\mathbb{R}^{n}$. Suppose that $\mathbf{u}$ is a random variable taking values in $\mathcal{U}$ with a smooth probability density that is strictly positive a.e.. If the pushforward $\mathbf{v}=\mathbf{h}(\mathbf{u})$ is uniformly distributed on $\mathcal{V}$, then $\mathbf{h}$ is a global diffeomorphism; in particular, $\mathbf{h}$ is bijective and depends on every component of $\mathbf{u}$ [^5] We do not claim originality for this result due to its fundamental nature in topology and measure theory; rather, we detail it here as a tool for our subsequent arguments.
@@
Let $p_{\mathbf{u}}:\mathcal{U}\to\mathbb{R}$ and $p_{\mathbf{v}}:\mathcal{V}\to\mathbb{R}$ denote the probability density functions of $\mathbf{u}$ and $\mathbf{v}$, respectively. Since $\mathbf{v}=\mathbf{h}(\mathbf{u})$, the change-of-variables formula (refer to, e.g., Bogachev, 2007) yields
@@
or equivalently, $\det J(\mathbf{h})(\mathbf{u})\neq 0$ a.e.. By the Inverse Function Theorem (see, e.g., Rudin, 1976), this implies that $\mathbf{h}$ is a local diffeomorphism.
Moreover, since $\mathcal{U}$ and $\mathcal{V}$ are simply connected, open $\mathcal{C}^{1}$ manifolds, standard covering space theory (refer to, e.g., the discussion around Theorem 1.38 in Hatcher, 2002) implies that $\mathbf{h}$ is a covering map. The uniformity of $p_{\mathbf{v}}$ forces $\mathbf{h}$ to be surjective (otherwise, some points in $\mathcal{V}$ would have zero density, contradicting uniformity). Since any covering map from a simply connected space is trivial, $\mathbf{h}$ must be equivalent to the identity covering. In other words, $\mathbf{h}$ is a homeomorphism onto $\mathcal{V}$ and hence both injective and surjective (i.e., a bijection).
Finally, the fact that the Jacobian determinant is nonzero a.e. guarantees that $\mathbf{h}$ depends on all components of $\mathbf{u}$; if any component were omitted, the rank of the Jacobian would drop, contradicting non-singularity. Furthermore, by the Global Inverse Function Theorem (refer to, e.g., Ruzhansky and Sugimoto, 2015), the inverse of $\mathbf{h}$ is smooth.
@@
    First, we show that the objective function $\mathcal{L}_{\text{SymAlignMaxEnt}}(\mathbf{f}_{x},\mathbf{f}_{t})$ achieves a global minimum value of 00. At this minimum, any pair of smooth functions $\mathbf{f}_{x}$ and $\mathbf{f}_{t}$ satisfying this condition must exhibit invariance across modalities. This invariance condition ensures that the learned image representations and text representations must align across all positive $\mathbf{x}$ and $\mathbf{t}^{\scriptscriptstyle(\theta)}$ pairs.
@@
    Next, we prove that minimizing $\mathcal{L}_{\text{SymAlignMaxEnt}}$ inherently eliminates any dependence of the learned representations on modality-specific variables $\mathbf{m}_{x}$ or $\mathbf{m}_{t}$. This ensures that the representations are restricted to the dependence on latent semantic variables.

### v4

@@ 5.2 MPI3D-Complex Dataset
The MPI3D-Complex dataset [18] contains real-world images annotated with seven mutually independent, discrete ground-truth factors: object color (color), shape (shape), size (size), camera height (cam.), background color (back.), horizontal position (hori.), and vertical position (vert.). We treat hori. and vert. as image-specific factors, and the remaining five as semantic variables. See Figure 5 for example images from the MPI3D-Complex dataset.
@@
Training is performed using the loss $\mathcal{L}_{\text{\tiny MMCL}}$ defined in Eq. 1, following the formulation in [12]. Performance is evaluated using the average Matthews Correlation Coefficient (MCC) across three random seeds, based on the prediction of all image latent variables from the learned image and text representations using linear and nonlinear MLP decoders, respectively. Full details on dataset attributes, text generation, bias configurations, and architectures are provided in § D.1.
@@ Results.
Figure 6 presents the nonlinear MCC results with learned representations. The findings demonstrate that, even with discrete latent variables, misaligned semantic variables across modalities, whether due to selection or perturbation biases, are consistently excluded from the representations (MCC $=0$). In contrast, unbiased semantics are well recovered, with MCC scores predominantly approaching 1 and all values $\geq 0.8$, reinforcing our theoretical findings. Further investigations into MCC using linear classifiers and ablation studies on encoder dimensionality are provided in § D.2.
@@ 5.3 Causal3DIdent Dataset
We further conduct experiments on the Causal3DIdent dataset [80, 66, 12, 76], a semi-synthetic dataset allowing explicitly defined causal relations among latent variables. Images are generated from 10 latent variables: 3 discrete (shape, x_pos, y_pos) and 7 continuous (color, s_pos, s_color, b_color, alpha, beta, gamma). We treat the rotation angles (alpha, beta, gamma) as image-specific, and the remaining as semantic variables following a causal graph among latent variables (Figure 15). See Figure 7 for example images from the dataset.
@@ Results.
Figure 8 presents the prediction performance of a nonlinear MLP classifier or regressor trained on learned image representations. We observe that unbiased semantic variables shared across modalities, whether continuous (e.g., s_pos) or discrete (e.g., shape, x_pos, y_pos), are reliably recovered by MMCL across all settings, with predictive performance approaching perfect ($R^{2}\approx 1$). For semantic variables that are continuous in the image modality but discretized in the text modality, prediction performance shows some degradation (e.g., s_color in selection setting 6 or perturbation setting 2), yet still achieves relatively high $R^{2}$ scores. Image-specific variables are consistently excluded from the learned representations, as indicated by $R^{2}=0$.
@@ 6 Related Work
Recent work has sought to formalize the theoretical foundations of multimodal and multi-view representation learning, particularly under contrastive objectives. Wang et al. [67] decompose the InfoNCE loss [49] into an alignment term, which pulls positive pairs together, and a uniformity term, which encourages dispersion over a hypersphere—laying the groundwork for subsequent analysis. Zimmermann et al. [80] show that contrastive objectives can invert the data-generating process, while Liu et al. [43] extend this result to multimodal settings. However, these approaches often rely on strong assumptions about latent distributions or manifold structures, which limits their practical applicability. A complementary line of work, such as [66], demonstrates that contrastive learning with data augmentation can recover shared content without requiring such strong parametric assumptions. This has been extended to MMCL by [46, 12], and Yao et al. [76] further investigate identifiability under partial observability in multi-view settings. Distinct from prior work, we do not assume fixed content-style decompositions, causal directions, or perfectly aligned pairs. Instead, we analyze MMCL with image-text pairs under cross-modal misalignment and systematically examine its impact on representation learning.
@@
Multimodal contrastive learning (MMCL) has achieved significant empirical success, particularly in aligning visual and textual modalities using models such as CLIP [53] and ALIGN [25]. These successes are partly attributed to the use of massive training corpora, e.g., LAION-5B [59], which are substantially larger than those used for vision-only foundation models [21, 50]. However, real-world multimodal datasets are often imperfectly aligned and noisy [47]. Existing empirical methods [1, 32] typically treat such misalignment as label noise, employing strategies such as multiple-instance learning or dataset refinement [1] to mitigate its impact. While some recent work suggests that contrastive models are robust to certain forms of structured misalignment [48], Others suggest augmenting text as a proxy for changes in visual content [5]. Our work suggests that cross-modal misalignment can act as either a barrier or a bridge, depending on the application. Unlike [48], which assumes linear representations without modeling the generative process, our analysis is grounded in a realistic latent variable model that provides a deeper understanding of misalignment from a data-generating perspective.
@@
Identifiability analysis addresses the fundamental question of whether the learning process can uniquely recover the latent generative structure or distribution underlying the observed data. This problem has been extensively studied in the context of nonlinear independent component analysis (ICA) [22, 44, 26, 61] and causal representation learning (CRL) [39, 41, 77, 75, 42]. In either setting, full identifiability—typically up to permutation—is rarely achievable in practice without strong assumptions, such as access to interventional data. Consequently, recent works have focused on partial identifiability [19, 29, 17] or relaxed equivalence classes, such as identifiability up to linear transformations [80, 43] or up to group-wise/block-wise indeterminacy [66, 76], which can offer sufficient guarantees for specific tasks or applications. In the context of multimodal representation learning, several recent studies have explored identifiability results [46, 12, 43], but largely neglect the presence of cross-modal misalignment. In contrast, our work explicitly models misalignment and adopts a block-identifiability definition to characterize the extent to which semantic factors can be recovered up to an invertible mapping.
@@
Invariant representation learning (IRL) seeks to learn representations that remain robust under distributional shifts between environments [2, 14, 17], particularly in settings where empirical risk minimization (ERM) [64] fails to generalize out of distribution. In the absence of such challenge, ERM is sufficient for in-distribution prediction, rendering the objective of IRL unnecessary. From a causal perspective, learning invariant representations—or more ambitiously, autonomous mechanisms [58]—requires observing sufficient variations in the latent factors underlying data [39]. Such variability can be introduced through interventional data [37, 31], exchangeability assumptions [54], or the use of auxiliary variables such as domain indices [77, 40]. However, direct interventions on latent variables are typically infeasible in real-world observational data, and auxiliary variable methods often require access to a large number of diverse environments to ensure identifiability—an assumption that is often challenging to satisfy in practice. Our work sheds light on an alternative approach: by leveraging the inherent flexibility of text supervision, we show that manipulating biases, specifically through selective omission or semantic perturbation in text, can serve as a controllable proxy for environmental variation.
@@ 7 Conclusion and Discussion
In this work, we present a formal analysis of cross-modal misalignment in multimodal data pairs within the MMCL framework, examining its impact on the learned representations. We demonstrate that contrastive multimodal encoders retain only the unbiased shared semantic variables, while discarding misaligned latent variables. When image-text pairs exhibit misalignment due to selection or perturbation biases, the joint embedding prioritizes consistent content, while omitting altered or missing aspects. This trade-off is fundamental: perfectly aligned text captions preserve rich semantic detail, whereas selective or biased text can enhance domain invariance by filtering out distribution-sensitive factors. Our experiments, conducted across simulations and image-text datasets, empirically validate these theoretical findings. These insights highlight the need for multimodal learning frameworks that mitigate misalignment or leverage beneficial biases to enhance representation learning in real-world settings. More discussion of these implications is provided in App. G.
@@ Appendix A Notation and Terminology
$\mathbb{I}_{\mathbf{s}}$ Index set of semantic variables: $[n_{s}]$
@@
$\mathcal{L}_{\text{\tiny MMCL}}(\mathbf{f}_{x},\mathbf{f}_{t})$ Symmetric InfoNCE loss for MMCL (Eq. 1)
$\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}(\mathbf{f}_{x},\mathbf{f}_{t})$ Alignment and entropy maximization loss (Eq. 2)
Notations for Cross-Modal Misalignment
$\theta$ Selection bias, an integer realization in the range $[2^{n_{s}}-1]$
@@
$\rho$ Perturbation bias, an integer realization in the range $[2^{|\mathbb{I}_{\theta}|}-1]$
@@
$\mathbf{s}_{\mathbb{I}^{c}_{\theta}}$ Omitted latent semantic variables when generating text
@@
$\mathbf{z}_{x}$ Combined latent variables for images: $(\mathbf{s},\mathbf{m}_{x})$
$\mathbf{z}_{t^{\scriptscriptstyle(\theta)}}$ Combined latent variables for text under $\theta$: $(\mathbf{s}_{\mathbb{I}_{\theta}},\mathbf{m}_{t})$
@@ B.1 Lemmas
Lemma B.1 (Global Minimum of $\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}$).
In the context of the proposed LVM as outlined in § 3.1, and Asms. 3.1 and 3.2, the global minimum of
$\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}(\mathbf{f}_{x},\mathbf{f}_{t})=\mathop{\mathbb{E}}_{(\mathbf{x},\mathbf{t}^{\scriptscriptstyle(\theta)})\sim p_{\mathbf{x},\mathbf{t}^{\scriptscriptstyle(\theta)}}}\left[\|\mathbf{f}_{x}(\mathbf{x})-\mathbf{f}_{t}(\mathbf{t}^{\scriptscriptstyle(\theta)})\|_{2}\right]-\frac{1}{2}\Bigl{(}H\bigl{(}\mathbf{f}_{x}(\mathbf{x})\bigr{)}+H\bigl{(}\mathbf{f}_{t}(\mathbf{t}^{\scriptscriptstyle(\theta)})\bigr{)}\Bigr{)},$ (6)
@@
  • The operator $(\cdot)_{\mathbb{I}^{c}_{\rho}}$ extracts the components corresponding to the unbiased semantic variables (i.e., unaffected by the selection bias nor the perturbation bias), with $n=\left|\mathbb{I}^{c}_{\rho}\right|$ being their dimensionality.
  • $\mathbf{d}=(d_{1},\dots,d_{n})$ is defined via the Darmois construction [11, 22, 66], where for each $i\in[n]$,
@@
We prove that the candidate functions $\mathbf{f}_{x}^{*}$ and $\mathbf{f}_{t}^{*}$ in Equations 7 and 8 yield $\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}(\mathbf{f}_{x}^{*},\mathbf{f}_{t}^{*})=0$. Substituting these candidate functions into the loss in Eq. 6, we have
$\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}(\mathbf{f}_{x}^{*},\mathbf{f}_{t}^{*})$ $=\mathop{\mathbb{E}}_{(\mathbf{x},\mathbf{t}^{\scriptscriptstyle(\theta)})\sim p_{\mathbf{x},\mathbf{t}^{\scriptscriptstyle(\theta)}}}\Bigl{[}\|\mathbf{f}_{x}^{*}(\mathbf{x})-\mathbf{f}_{t}^{*}(\mathbf{t}^{\scriptscriptstyle(\theta)})\|_{2}\Bigr{]}$
@@
$\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}(\mathbf{f}_{x}^{*},\mathbf{f}_{t}^{*})$ $=\mathop{\mathbb{E}}_{(\mathbf{z}_{x},\mathbf{z}_{t^{\scriptscriptstyle(\theta)}})\sim p_{\mathbf{z}_{x},\mathbf{z}_{t^{\scriptscriptstyle(\theta)}}}}\Bigl{[}\|\mathbf{d}(\mathbf{s}_{\mathbb{I}^{c}_{\rho}})-\mathbf{d}(\tilde{\mathbf{s}}_{\mathbb{I}^{c}_{\rho}})\|_{2}\Bigr{]}$
@@
$\tilde{\mathbf{s}}_{\mathbb{I}^{c}_{\rho}}=\mathbf{s}_{\mathbb{I}^{c}_{\rho}}\quad\text{almost surely (a.s.)}\quad\forall\,\mathbf{s}\sim p_{\mathbf{s}},\tilde{\mathbf{s}}_{\mathbb{I}_{\theta}}\sim p_{\tilde{\mathbf{s}}_{\mathbb{I}_{\theta}}\mid\mathbf{s},\theta,\rho},$
@@
Further, by the properties of the Darmois construction [11], the mapping $\mathbf{d}$ transforms $\mathbf{s}_{\mathbb{I}^{c}_{\rho}}$ into a uniform distribution over $(0,1)^{n}$ (with $n=|\mathbb{I}^{c}_{\rho}|$). Since the uniform distribution is the unique maximum entropy (i.e., zero) distribution on a bounded domain (under no further moment constraints) [24, 9], the entropy terms in the loss are maximized. In the formulation of $\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}$, this maximal entropy precisely cancels any potential reduction in the loss, ensuring that
$\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}(\mathbf{f}_{x}^{*},\mathbf{f}_{t}^{*})=0.$
Therefore, the global minimum of $\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}$ is achieved at 0 by the given function pairs $\mathbf{f}_{x}^{*}$ and $\mathbf{f}_{t}^{*}$, completing the proof. ∎
@@
Let $\mathbf{h}:\mathcal{U}\to\mathcal{V}$ be a smooth map between simply connected, open $\mathcal{C}^{1}$ manifolds $\mathcal{U},\mathcal{V}\subseteq\mathbb{R}^{n}$. Suppose that $\mathbf{u}$ is a random variable taking values in $\mathcal{U}$ with a smooth probability density that is strictly positive a.e.. If the pushforward $\mathbf{v}=\mathbf{h}(\mathbf{u})$ is uniformly distributed on $\mathcal{V}$, then $\mathbf{h}$ is a global diffeomorphism; in other words, $\mathbf{h}$ is bijective and depends on every component of $\mathbf{u}$ [^7] We do not claim originality for this result due to its fundamental nature in topology and measure theory; rather, we detail it here as a tool for our subsequent arguments.
@@
Let $p_{\mathbf{u}}:\mathcal{U}\to\mathbb{R}$ and $p_{\mathbf{v}}:\mathcal{V}\to\mathbb{R}$ denote the probability density functions of $\mathbf{u}$ and $\mathbf{v}$, respectively. Since $\mathbf{v}=\mathbf{h}(\mathbf{u})$, the change-of-variables formula (refer to, e.g., [4]) yields
@@
or equivalently, $\det J(\mathbf{h})(\mathbf{u})\neq 0$ a.e.. By the Inverse Function Theorem (see, e.g., [56]), this implies that $\mathbf{h}$ is a local diffeomorphism.
Moreover, since $\mathcal{U}$ and $\mathcal{V}$ are simply connected, open $\mathcal{C}^{1}$ manifolds, standard covering space theory (refer to, e.g., the discussion around Theorem 1.38 in [20]) implies that $\mathbf{h}$ is a covering map. The uniformity of $p_{\mathbf{v}}$ forces $\mathbf{h}$ to be surjective (otherwise, some points in $\mathcal{V}$ would have zero density, contradicting uniformity). Since any covering map from a simply connected space is trivial, $\mathbf{h}$ must be equivalent to the identity covering. In other words, $\mathbf{h}$ is a homeomorphism onto $\mathcal{V}$ and hence both injective and surjective (i.e., a bijection).
Finally, the fact that the Jacobian determinant is nonzero a.e. guarantees that $\mathbf{h}$ depends on all components of $\mathbf{u}$; if any component were omitted, the rank of the Jacobian would drop, contradicting non-singularity. Furthermore, by the Global Inverse Function Theorem (refer to, e.g., [57]), the inverse of $\mathbf{h}$ is smooth.
@@
    First, we show that the objective function $\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}(\mathbf{f}_{x},\mathbf{f}_{t})$ achieves a global minimum value of 00. At this minimum, any pair of smooth functions $\mathbf{f}_{x}$ and $\mathbf{f}_{t}$ satisfying this condition must exhibit invariance across modalities. This invariance condition ensures that the learned image representations and text representations must align across all positive $\mathbf{x}$ and $\mathbf{t}^{\scriptscriptstyle(\theta)}$ pairs.
@@
    Next, we prove that minimizing $\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}$ inherently eliminates any dependence of the learned representations on modality-specific variables $\mathbf{m}_{x}$ or $\mathbf{m}_{t}$. This ensures that the representations are restricted to the dependence on latent semantic variables.